1/2

Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,
El Hechicero  The Werewolf  Varulven
übersetzt von Hans Eiseneck  übersetzt von Alexander Gross  übersetzt von Sven Collberg

Un hechicero, con esposa vino un día a la fosa de su maestro, y al profesor pidió: “¿Me cuentas? ¡Por favor!”

A Werewolf, troubled by his name, Left wife and brood one night and came To a hidden graveyard to enlist The aid of a long-dead philologist.

En varulv lopp från fru m.m. en vacker afton samt begav sig till en bymagisters grav och bad: “Var snäll och konjugera.

El profesor sin vacilar subió y comenzó a hablar. Desde allá de su letrero se dirigió al hechicero:

“Oh sage, wake up, please don’t berate me,” He howled sadly, “Just conjugate me.” The seer arose a bit unsteady Yawned twice, wheezed once, and then was ready. mig!” Byskolläraren stod upp på blecknamnsskyltens mässingknopp med svar till den, som satt på pass med tåligt korslagd tass på tass.

“El hechiuno” le explica, “el hechidos” después indica, “el hechitres” dice al fin, “con eso basta ¡chiquitín!”

“Well, ‘Werewolf’ is your plural past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present tense, And ‘Iswolf,’ ‘Arewolf’ in this same sense.”

“Jag var-ulv 1:a pers. i sing., du, han, hon, den, det var-ulv. Så vi voro-ulv, I voren. På de voro-ulv kom ingenting.”

El hechicero, conmovido por lo que capta su oído, al maestro ruega, encantado, que le adjunte un quebrado.

“I know that–I’m no mental cripple– The future form and participle Are what I crave,” the beast replied. The scholar paused–again he tried:

Varulven myste gladåt hur

pers. hade följts av sing. och plur. “Fast”, bad han, “kan det inte ges ens en enda pers. i tempus presens?”

Pero el profesor confiesa que nada sabe con certeza: “Hay muchos hechos. Hechiceros únicamente hay enteros.”

“A ‘Will-be-wolf?’ It’s just too long: ‘Shall-be-wolf?’ ‘Has-been-wolf?’ Utterly wrong! Such words are wounds beyond all suture– I’m sorry, but you have no future.”

Bekänna måste vår magister, att därvidlag fanns vissa brister. En varulv sine imperfecto var contradictio in adjecto.

…  …  …

2/2

El brujo se volvió muy triste: “Mi cara mitad, ¿no la viste?” Ya que no era sabio con gratitud se despidió.

The Werewolf knew better–his sons still slept At home, and homewards now he crept, Happy, humble, without apology For such folly of philology.

Varulven satt med tårad lins.

Han visste att han fanns och finns.

Därtill kom också fru m.m. som även måste existera.

Übersetzung des deutschen Gedichtes “Der Werwolf” von Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) ins hispana durch Hans Eiseneck.

Arg-825-1656 (2013-02-13 18:46:00)

Tiu ĉi trsduko en la kastilian lingvon troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Übersetzung des deutschen Gedichtes “Der Werwolf” von Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) ins Englische durch Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Übersetzung des deutschen Gedichtes “Der Werwolf” von Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) ins Schwedische durch Sven Collberg.

Arg-825-1654 (2013-02-13 17:00:43)

Tiu ĉi francigo troviĝas ne la reterjo http: // bertilow. com/ literaturo/ lupfantomo. html# sv .

http://www.poezio.net/version?poem-id=825&version-id=1656,1657,1654